(this is a word problem involving quadratic equation)
an object is projected directly upward from the ground. after t seconds its distance in feet above the ground is s=144t-16r^2 after how many sec. will the object be 128 feet above the ground? (please explain how u got the answer and use the factoring method)

Answer by stanbon(75887)

an object is projected directly upward from the ground. after t seconds its distance in feet above the ground is s=144t-16r^2 after how many sec. will the object be 128 feet above the ground?
--------------
When the object is on the ground, s=zero.
---
Solve: -16t^2+144t = 0
-16t(t-9) = 0
t = 0 seconds or t = 9 seconds
----
That means the object was on the ground with
time started and is back on the ground after
9 seconds.
=============
Cheers,
Stan H.